1. Solve the given differential equation by undetermined coefficients.
y'' + 6y' + 9y = −xe^4x

y(x) =____

2. Solve the given differential equation by undetermined coefficients.
y''' − 3y'' + 3y' − y = e^x − x + 21

y(x)= _____
1
Expert's answer
2020-07-09T20:08:21-0400

1.The given equation can be written as, "(D^{2}+6D+9)y = -xe^{4x}." The auxiliary equation is,

"m^2+6m+9=0." Solving we get, "m=-3(twice)". The complementary function is,

"y_{c} = (c_{1}+c_{2}x)e^{-3x}" . Let "y_{p} = (Ax+B)e^{4x}" be the trial particular solution.

Then, "y_{p}" satisfies the given differential equation. Therefore, "y''_{p}+6y'_{p}+9y_{p} = -xe^{4x}" .

Now,

"\begin{aligned}
y'_{p} &= e^{4x} \cdot A+ (Ax+B)\cdot e^{4x} \cdot 4\\ &= e^{4x}(4Ax+A+4B)\\
y''_{p}&= e^{4x} \cdot 4A + (4Ax+A+4B)\cdot e^{4x}\cdot 4\\
&= e^{4x}(16Ax+8A+16B)
\end{aligned}"

Hence,

"\begin{aligned}
e^{4x}\bigg(16Ax+8A+16B+6\cdot(4Ax+A+4B)+9 \cdot (Ax+B)\bigg) &= -xe^{4x}\\
49Ax+14A+49B &= -x\\
\end{aligned}"

Comparing the coefficients and constant terms and solving we get,

"A = -\dfrac{1}{49}, B = \dfrac{2}{343}" . Therefore, "y_{p} = \left(-\dfrac{x}{49}+\dfrac{2}{343}\right)e^{4x} = -\dfrac{e^{4x}}{49}\left(x-\dfrac{2}{7}\right)" .

The general solution is,

"y = (c_{1}+c_{2}x)e^{-3x} - \dfrac{e^{4x}}{49}\left(x-\dfrac{2}{7}\right)"


2.The given equation can be written as, "(D^{3}-3D^{2}+3D-1)y=e^{x}-x+21". The auxiliary equation is "m^{3}-3m^{2}+3m-1=0." Solving we get, "m = 1 (thrice)." The complementary function is, "y_{c} = (c_{1}+c_{2}x+c_{3}x^{2})e^{x}." Since the complementary function is a linear combination of the term "e^{x}", we choose the trial solution as "y_{p} = Ax^{3}e^{x} + Bx + C."

Thus,

"\begin{aligned}
y'_{p} &= A(x^{3}e^{x}+3x^{2}e^{x})+B = Ae^{x}(x^{3}+3x^{2})+B\\
y''_{p} &= A\left(e^{x}(3x^{2}+6x)+e^{x}(x^{3}+3x^{2})\right)\\
&= Ae^{x}\left(x^{3}+6x^{2}+6x\right)\\
y'''_{p} &= A\left(e^{x}(3x^{2}+12x+6)+e^{x}(x^{3}+6x^{2}+6x)\right)\\
&=Ae^{x}(x^{3}+9x^{2}+18x+6)
\end{aligned}"

Using this in the given differential equation we get,

"\begin{aligned}
e^{x}\bigg(Ax^{3}+9Ax^{2}+18Ax+6A\\ -3\cdot(Ax^{3}+6Ax^{2}+6Ax) \\ +3\cdot(Ax^{3}+3Ax^{2})-Ax^{3}\bigg) \\ +3B - Bx - C &= e^{x}-x+21\\
6Ae^{x}-Bx+(3B-C)&=e^{x}-x+21
\end{aligned}"

Comparing the coefficients, constant terms and solving we get,

"A=\dfrac{1}{6}, B = 1, C = 18."


Hence, "y_{p}= \dfrac{x^{3}e^{x}}{6}+x-18."

The general solution is,

"y=(c_{1}+c_{2}x+c_{3}x^{2})e^{x}+\dfrac{x^{3}e^{x}}{6}+x-18."